Home‎ > ‎Reports‎ > ‎Crew097‎ > ‎

day13


Crew 97 Log Book for January 13, 2011

Photos of the day





Commander's Check-In Report

Judah Epstein Reporting

Crew Physical Status:
Good.

Time Departed/Returned from EVA:
EVA #27: 09:40 / 11:55. Jim, Nathan. EVA #28: 14:05 / 15:09. Amanda, Nathan, Jim.

Brief Narrative of Field Mission Results:
EVAs today consisted of Rock Varnish sample collecting, Water Gully analysis, and Satellite data comparison collection.

EVA Data/Interpretations:
EVA #27: Satellite comparison data was collected. EVA #28: Rock Varnish samples were collected and Water Gullies were analyzed.

Engineering/Hab Maintenance:
Top layer of clogged sand removed from greywater straining tank.

Report Transmission Schedule:
CMDR Check-in, ENGR, CMDR Report, Photos, 2 EVA Reports, 3 Science Reports

Plans for Tomorrow:
- Cooking/Cleaning - Amanda/Nathan
- 07:30 - Breakfast
- 8:00 - Exercise
- 8:20-9:30 - Crew-member psychophysiological testing
- Engineering checks
- EVA #29: Satellite - Jim, Amanda
- 11:30 Lunch
- 12:00-13:10 - Crew-member psychophysiological testing
- EVA #30: Satellite / Varnish - Nathan, Jim, Judah
- 15:00 - Finish Sim
- Dinner
- Crew meeting about Saturday and handover
- Psychophysiological crew analysis
- Reports, Food Study
- Work on experiments
- Crew-member Imaginiff game
- Showers - Lucinda/Tonya

Inventory:
n/a

Support Requested:
n/a

Miscellaneous:
n/a


Commander's Report

Judah Epstein Reporting

Today was Crew 97's last full day in Sim. The team continued scientific experiments of psychophysiological testing and water filtration analysis. EVAs today consisted of Rock Varnish collection to test for signs of life, satellite data comparison, and Water Gully analysis.

The crew received their initial Alpine Aire dehydrated foods and have begun to enjoy the new foods and ease of preparation. Additionally, the hot water heater is now functioning and therefore the crew have begun to take showers (whereas past shower opportunities were skipped because of the freezing cold water).

In the Greenhab, the top layer of clogged sand was removed from the strainer tank to increase flow rate of the greywater filtration system.

The crew is now preparing for their final day in Simulation with a morning and afternoon EVA before the scheduled 3:00pm return back to Earth.

Science Report 1

Judah Epstein Reporting

The purpose of this project is to explore the effectiveness of a slow sand sedimentary water filtration system in a planetary environment. To simulate a planetary environment the experiment is conducted at the Mars Desert Research Station, a self-sustaining planetary research experiment and simulation expedition.

Background information regarding the project abstract, purpose, goals, and specific experiment details and information are included in the Science Report from January 6.

Data collected from Jan 12 and prior is available in this spreadsheet.


Updates on results:

Filter 1 (Martian Simulant) and Filter 2 (Commercial Sand) are not achieving optimal results as the filter is clogged with extremely low hydraulic conductivity. Therefore not enough effluent water was produced to take measurements with the Qanta probe (Temp, SpC, DO, pH, TDS, DO%, ORP).

Turbidity results: Turbidity of effluent from all three slow sand filters has generally been increasing. Turbidity of effluent from the wetland, greywater, and Hab potable are remaining generally consistent.

Science Report 2

Tonya Thompson Reporting

I performed 3 psychophysiological reading sessions on each crewmember today. Each session lasted 3 minutes.

The crewmember entered my stateroom, sat in a chair, placed finger sensors on their 3 middle fingers, and watched a white image projected onto a wall for 3 minutes. During that time, I recorded their heart rate variability (HRV) and skin conductance level (SCL). The data for each session was saved to the hard drive of a desktop pc in my stateroom.

I also had each crewmember fill out a Cohen’s Perceived Stress Scale questionnaire at the end of the evening after dinner.

Engineering Report

Nathan Wong Reporting

Generator/Electricity:

Kitty(Diesel Generator) (Used? Yes/No): Yes

Honey (Gas Generator) (Used? Yes/No): No

Battery Charge Level:

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.50

Propane Tank (Full/0.75/0.5/0.25/Empty): 0.47

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 3 Gas cans full

Notes/Comments:


Oil:

Kitty Oil Quantity (# of Quarts in storage): 3

Honey Oil Quantity (# of Quarts in storage): 2

ATV Oil Quantity (# of Quarts in storage): 2

Car Oil Quantity (# of Quarts in storage): none


Water Status:

Outside Potable Water Tank Level (inches from bottom): 10%

Trailer Potable Water Tank Level (inches from bottom): 75%

Hab Potable Water Tank Level (inches from bottom): half

Potable Water Meter Reading: 40857.8

In to GreenHab Meter Reading:

Notes/Comments: Water line worked today.


GreenHab:

Condition of plants in Tank 1: None

Condition of plants in Tank 2: 3 plants

Condition of plants in Tank 3: 7 plants

Greenhab Notes/Comments:
We removed an inch of sand from the filter (the clogged top layer). Tomorrow we will send pictures of before and after of the water. Water quality analysis can also be viewed in the Science Report of the Water Filtration experiment; specifically "Filter 4: MDRS Hab Wetland Filtration".


Observatory:

Telescope Used? (Yes/No): No

Observatory Notes/Comments: Observatory not used


Transportation:

Hab Car (NOT RENTAL)
Used (yes/no): No
Oil Checked (yes/no): No
Tire Status: good
Odometer at end of day: 123085
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: Yes
Oil Checked: No
Fuel Consumed: Yes
Tire Status: Inflated
Comments:

ATV 2 (Spirit):
Used: Yes
Oil Checked: No
Fuel Consumed: Yes
Tire Status: Inflated
Comments:

ATV 3 (Viking I):
Used: No
Oil Checked: No
Fuel Consumed: No
Tire Status: Inflated
Comments:

ATV 4 (Viking II):
Used: Yes
Oil Checked: No
Fuel Consumed: Yes
Tire Status: Inflated
Comments:


Heating and Ventilation:
Thermostat setting upstairs (Farenheit): off
Actual temperature upstairs (Farenheit): 75

Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): 57


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: All webcams are now functioning properly.


General Engineering Notes/Comments:

-Sand was removed from the filter tank.

-A heater was placed on the toilet line to see if it was frozen

-The line leading from the Hab to the grey water tank is leaking.

EVA 27 Report

Jim Crowell Reporting

Time: 09:40 / 11:55

Crew Members: Jim Crowell, Nathan Wong

Site Location: South of the Hab

Transit Mode: ATVs / Walking

Objectives: To take two panoramas south of the Hab

Results: Both desired locations reached and panoramas taken.

Lessons Learned: Even if failed on the first attempt, by looking for an alternative, one can accomplish what is needed. One of the pans needed was at the top of a mountain. After the first attempt to reach the top failed, another way up was sought out and found.


EVA 28 Report

Amanda Damptz Reporting

Time: 2:05 pm to 3:09 pm

Crew Members: Jim Crowell, Nathan Wong and Amanda Damptz

Site Location: North East of the Hab

Transit Mode: ATV

Objectives: Take photographs of gullies and collect rock varnish samples.

Results: Three photographs of gullies taken. GPS coordinates taken. Two samples of rock varnish collected.

Lesson Learned: The wrong type of rock hammer was taken out delaying the sample collection process.